# Uncertainty Calibration for Counterfactual Propensity Estimation in Recommendation

### v2

@@ IV-C CVR Prediction Results of IPS(RQ2)
We utilize both uncalibrated and calibrated propensity scores to train the debiasing CVR models $f_{\theta}$, respectively. As a baseline, we train the recommendation model without using propensity scores, implying that all samples are given equal weight for loss. Table II presents the overall IPS debiasing performance in terms of DCG@K and Recall@K ($K=2,4,6$) on three real-world datasets [^4] Recall refers to the recall number, which may exceed 1 based on the equation in Eqn. 25. We repeat the experiments ten times and report the mean results to mitigate randomness. From the table, it can be observed that the IPS method with uncalibrated propensity scores achieves marginal improvement in recommendation performance compared to the baseline methods. Interestingly, the recall metric for the Yahoo! R3 dataset even shows a slight decrease, indicating that poorly calibrated propensity scores do not effectively aid the IPS-based training process.
With calibrated propensity scores, the IPS debiasing method shows significant improvement. As demonstrated in Table II and Table V, Platt scaling calibrated propensity scores outperform the uncalibrated ones in terms of DCG@K and Recall@K ($K=2,4,6$) on three real-world datasets. Specifically, Platt scaling shows a substantial relative improvement of 2.08%, 1.98%, and 2.07% over the uncalibrated method for DCG@2, DCG@4, and DCG@6, respectively.
@@ IV-D CVR predcition Results of SOTA debiasing methods(RQ3)
As our method improves the quality of propensity score estimation, it can readily be extended to other propensity score-based debiasing methods. We conducted experiments on six state-of-the-art CVR prediction models: DR-JL [10], MRDR [1], GPL [4], CDR [46], DR-V2 [48], and ESCM2 [3]. Table IV demonstrates that calibrated propensity scores outperform raw propensity scores on all evaluation metrics, highlighting the effectiveness of uncertainty calibration for other propensity score-based methods. Table VI shows that our method surpasses current state-of-the-art (SOTA) approaches in terms of performance.
@@ IV-E Efficiency experiment
DCG@K Recall@K
Baseline Dataset Methods K=2 K=4 K=6 K=2 K=4 K=6 Average
Raw 0.7454$\pm$0.0158 1.0185$\pm$0.0196 1.2014$\pm$0.0126 0.8717$\pm$0.0214 1.4570$\pm$0.0243 1.9489$\pm$0.0209 1.2071
Dropout 0.7496$\pm$0.0205 1.0271$\pm$0.0165 1.2073$\pm$0.0157 0.8835$\pm$0.0249 1.4764$\pm$0.0216 1.9603$\pm$0.0286 1.2173
Ensembles 0.7546$\pm$0.0160 1.0254$\pm$0.0110 1.2132$\pm$0.0119 0.8835$\pm$0.0228 1.4637$\pm$0.0126 1.9684$\pm$0.0232 1.2181
Coat Shopping Platt 0.7539$\pm$0.0201 1.0389$\pm$0.0208 1.2241$\pm$0.0170 0.8852$\pm$0.0235 1.4940$\pm$0.0290 1.9928$\pm$0.0208 1.2315
Raw 0.5450$\pm$0.0093 0.7405$\pm$0.0097 0.8779$\pm$0.0091 0.6402$\pm$0.0113 1.0587$\pm$0.0143 1.4112$\pm$0.0141 0.8789
Dropout 0.5501$\pm$0.0152 0.7597$\pm$0.0126 0.8774$\pm$0.0103 0.6539$\pm$0.0161 1.0822$\pm$0.0124 1.4183$\pm$0.0065 0.8902
Ensembles 0.5431$\pm$0.0053 0.7491$\pm$0.0043 0.8747$\pm$0.0050 0.6510$\pm$0.0044 1.0923$\pm$0.0054 1.4293$\pm$0.0068 0.8899
DR-JL Yahoo! R3 Platt 0.5532$\pm$0.0058 0.7555$\pm$0.0042 0.8816$\pm$0.0034 0.6602$\pm$0.0059 1.0926$\pm$0.0073 1.4314$\pm$0.0062 0.8957
Raw 0.6830$\pm$0.0151 0.9661$\pm$0.0131 1.1578$\pm$0.0101 0.8185$\pm$0.0163 1.4261$\pm$0.0192 1.9409$\pm$0.0205 1.1654
Dropout 0.7255$\pm$0.0200 0.9946$\pm$0.0138 1.1847$\pm$0.0127 0.8438$\pm$0.0257 1.4177$\pm$0.0247 1.9282$\pm$0.0312 1.1824
Ensembles 0.7292$\pm$0.0273 1.0001$\pm$0.0192 1.1846$\pm$0.0182 0.8523$\pm$0.0352 1.4303$\pm$0.0190 1.9282$\pm$0.0299 1.1875
Coat Shopping Platt 0.7648$\pm$0.0190 1.0155$\pm$0.0170 1.2223$\pm$0.0165 0.8987$\pm$0.0230 1.4688$\pm$0.0258 1.9957$\pm$0.0395 1.2276
Raw 0.5371$\pm$0.0447 0.7441$\pm$0.0502 0.8636$\pm$0.0455 0.6383$\pm$0.0508 1.0808$\pm$0.0638 1.4006$\pm$0.0524 0.8774
Dropout 0.5458$\pm$0.0194 0.7461$\pm$0.0240 0.8718$\pm$0.0225 0.6547$\pm$0.0209 1.0821$\pm$0.0363 1.4199$\pm$0.0382 0.8867
Ensembles 0.5410$\pm$0.0093 0.7443$\pm$0.0098 0.8731$\pm$0.0090 0.6444$\pm$0.0120 1.0809$\pm$0.0133 1.4256$\pm$0.0127 0.8847
MRDR Yahoo! R3 Platt 0.5623$\pm$0.0092 0.7571$\pm$0.0066 0.8858$\pm$0.0072 0.6687$\pm$0.0103 1.0862$\pm$0.0080 1.4326$\pm$0.009 0.8988
TABLE IV: DRJL and MRDR CVR prediction performance on Coat Shopping and Yahoo! R3. The best results are shown in boldface and the second best results are marked using underline.
DCG@K Recall@K
@@
Raw 0.4426$\pm$ 0.0076 0.6728$\pm$ 0.0110 0.8471$\pm$ 0.0108 0.5404$\pm$ 0.0096 1.0351$\pm$ 0.0166 1.5028$\pm$ 0.0163 0.8401
MC Dropout 0.4515$\pm$ 0.0063 0.6855$\pm$ 0.0064 0.8593$\pm$ 0.0069 0.5520$\pm$ 0.0073 1.0545$\pm$ 0.0080 1.5208$\pm$ 0.0101 0.8539
Deep Ensembles 0.4562$\pm$ 0.0045 0.6893$\pm$ 0.0048 0.8627$\pm$ 0.0042 0.5579$\pm$ 0.0052 1.0575$\pm$ 0.0063 1.5228$\pm$ 0.0058 0.8577
IPS Platt Scaling 0.4657$\pm$ 0.0030 0.7021$\pm$ 0.0025 0.8774$\pm$ 0.0026 0.5677$\pm$ 0.0029 1.0754$\pm$ 0.0030 1.5455$\pm$ 0.0036 0.8723
Raw 0.4442$\pm$ 0.0083 0.6742$\pm$ 0.0111 0.8481$\pm$ 0.0115 0.5420$\pm$ 0.0096 1.0362$\pm$ 0.0160 1.5026$\pm$ 0.0170 0.8412
MC Dropout 0.4504$\pm$ 0.0042 0.6839$\pm$ 0.0044 0.8580$\pm$ 0.0052 0.5504$\pm$ 0.0041 1.0520$\pm$ 0.0047 1.5189$\pm$ 0.0071 0.8523
Deep Ensembles 0.4524$\pm$ 0.0064 0.6854$\pm$ 0.0039 0.8606$\pm$ 0.0043 0.5528$\pm$ 0.0051 1.0530$\pm$ 0.0048 1.5231$\pm$ 0.0067 0.8546
DR-JL Platt Scaling 0.4701$\pm$ 0.0073 0.7070$\pm$ 0.0079 0.8834$\pm$ 0.0080 0.5733$\pm$ 0.0081 1.0823$\pm$ 0.0095 1.5553$\pm$ 0.0096 0.8786
Raw 0.4369$\pm$ 0.0246 0.6531$\pm$ 0.0287 0.8144$\pm$ 0.0295 0.5305$\pm$ 0.0292 0.9957$\pm$ 0.0376 1.4290$\pm$ 0.0401 0.8099
MC Dropout 0.4412$\pm$ 0.0125 0.6659$\pm$ 0.0183 0.8305$\pm$ 0.0200 0.5376$\pm$ 0.0152 1.0205$\pm$ 0.0275 1.4630$\pm$ 0.0302 0.8265
Deep Ensembles 0.4406$\pm$ 0.0101 0.6681$\pm$ 0.0127 0.8432$\pm$ 0.0135 0.5390$\pm$ 0.0131 1.0275$\pm$ 0.0171 1.4972$\pm$ 0.0215 0.8359
MRDR Platt Scaling 0.4928$\pm$ 0.0057 0.7259$\pm$ 0.0076 0.8963$\pm$ 0.0079 0.5971$\pm$ 0.0070 1.0972$\pm$ 0.0109 1.5549$\pm$ 0.0119 0.8940
TABLE V: Overall performance on KuaiRand.
DCG@K Recall@K
Datasets Methods K=2 K=4 K=6 K=2 K=4 K=6 Average
MRDR-GPL 0.7488$\pm$ 0.0201 1.0061$\pm$ 0.0222 1.1949$\pm$ 0.0243 0.8734$\pm$ 0.0287 1.4219$\pm$ 0.0301 1.9283$\pm$ 0.0405 1.1956
MRDR-CDR 0.7579$\pm$ 0.0201 1.0192$\pm$ 0.0192 1.1991$\pm$ 0.0198 0.8903$\pm$ 0.0204 1.4515$\pm$ 0.0277 1.9325$\pm$ 0.0402 1.2084
DR-V2 0.7746$\pm$ 0.0188 1.0116$\pm$ 0.0164 1.2076$\pm$ 0.0176 0.8841$\pm$ 0.0184 1.4568$\pm$ 0.0152 1.9494$\pm$ 0.0324 1.2140
ESCM2-DR 0.7528$\pm$ 0.0177 1.0273$\pm$ 0.0189 1.2081$\pm$ 0.0229 0.8945$\pm$ 0.0186 1.4810$\pm$ 0.0162 1.9662$\pm$ 0.0299 1.2217
Coat Shopping MRDR-CAL(Ours) 0.7648$\pm$ 0.0190 1.0155$\pm$ 0.0170 1.2223$\pm$ 0.0165 0.8987$\pm$ 0.0230 1.4688$\pm$ 0.0258 1.9957$\pm$ 0.0395 1.2276
MRDR-GPL 0.5384$\pm$ 0.0194 0.7369$\pm$ 0.0252 0.8605$\pm$ 0.0211 0.6408$\pm$ 0.0197 1.0657$\pm$ 0.0222 1.3982$\pm$ 0.0241 0.8734
MRDR-CDR 0.5417$\pm$ 0.0162 0.7456$\pm$ 0.0123 0.8698$\pm$ 0.0125 0.6490$\pm$ 0.0202 1.0842$\pm$ 0.0182 1.4175$\pm$ 0.0147 0.8846
DR-V2 0.5518$\pm$ 0.0125 0.7479$\pm$ 0.0143 0.8732$\pm$ 0.0156 0.658$\pm$ 0.0111 1.0784$\pm$ 0.0181 1.4154$\pm$ 0.0158 0.8875
ESCM2-DR 0.5541$\pm$ 0.0144 0.7502$\pm$ 0.0126 0.8771$\pm$ 0.0171 0.6564$\pm$ 0.0156 1.0772$\pm$ 0.0175 1.4183$\pm$ 0.0154 0.8889
Yahoo! R3 MRDR-CAL(Ours) 0.5623$\pm$ 0.0092 0.7571$\pm$ 0.0066 0.8858$\pm$ 0.0072 0.6687$\pm$ 0.0103 1.0862$\pm$ 0.0080 1.4326$\pm$ 0.0090 0.8988
MRDR-GPL 0.4335$\pm$ 0.0123 0.6446$\pm$ 0.0178 0.8049$\pm$ 0.0143 0.5261$\pm$ 0.0171 0.9792$\pm$ 0.0178 1.4101$\pm$ 0.0146 0.7997
MRDR-CDR 0.4326$\pm$ 0.0098 0.6573$\pm$ 0.0145 0.8297$\pm$ 0.0132 0.5289$\pm$ 0.0126 1.0111$\pm$ 0.0146 1.4739$\pm$ 0.0165 0.8223
DR-V2 0.4465$\pm$ 0.0078 0.6795$\pm$ 0.0121 0.8533$\pm$ 0.0098 0.5452$\pm$ 0.0100 1.0459$\pm$ 0.0122 1.5117$\pm$ 0.0123 0.8470
ESCM2-DR 0.4773$\pm$ 0.0101 0.7059$\pm$ 0.0155 0.8760$\pm$ 0.0121 0.5781$\pm$ 0.0143 1.0694$\pm$ 0.0143 1.5258$\pm$ 0.0152 0.8721
KuaiRand MRDR-CAL(Ours) 0.4928$\pm$ 0.0057 0.7259$\pm$ 0.0076 0.8963$\pm$ 0.0079 0.5971$\pm$ 0.0070 1.0972$\pm$ 0.0109 1.5549$\pm$ 0.0119 0.8940
TABLE VI: The comparison with the SOTA methods

### v3

@@ IV-C CVR Prediction Results of IPS(RQ2)
We utilize both uncalibrated and calibrated propensity scores to train the debiasing CVR models $f_{\theta}$, respectively. As a baseline, we train the recommendation model without using propensity scores, implying that all samples are given equal weight for loss. Table II presents the overall IPS debiasing performance in terms of DCG@K, Recall@K ($K=2,4,6$) and AUC on three real-world datasets [^4] Recall refers to the recall number, which may exceed $1$.. We repeat the experiments ten times and report the mean results to mitigate randomness. From the table, it can be observed that the IPS method with uncalibrated propensity scores achieves marginal improvement in recommendation performance compared to the baseline methods. Interestingly, the recall metric for the Yahoo! R3 dataset even shows a slight decrease, indicating that poorly calibrated propensity scores do not effectively aid the IPS-based training process.
With calibrated propensity scores, the IPS debiasing method shows significant improvement. As demonstrated in Table II and Table V, Platt scaling calibrated propensity scores outperform the uncalibrated ones in terms of all evaluation metrics on three real-world datasets. For instance, Platt scaling based IPS demonstrates substantial relative improvements of 1.51%, 2.08%, 1.98%, and 2.07% over the uncalibrated IPS method for AUC, DCG@2, DCG@4, and DCG@6 on the Coat Shopping dataset, respectively.
@@ IV-D CVR predcition Results of SOTA debiasing methods(RQ3)
As our method improves the quality of propensity score estimation, it can readily be extended to other propensity score-based debiasing methods. We conducted experiments on six state-of-the-art CVR prediction models: DR-JL [10], MRDR [1], GPL [4], CDR [51], DR-V2 [53], and ESCM2 [3]. Table IV demonstrates that calibrated propensity scores outperform raw propensity scores on all evaluation metrics, highlighting the effectiveness of uncertainty calibration for other propensity score-based methods. Table VI shows that our method surpasses current state-of-the-art (SOTA) approaches in terms of performance. Table VI demonstrates that our method outperforms the current state-of-the-art (SOTA) approaches in terms of performance. This improvement is largely attributed to our method’s ability, particularly when combined with Platt scaling calibration, to more accurately estimate the probability of each (user, item) pair appearing in the observed data. Enhanced accuracy in propensity score estimation directly translates to higher-quality recommendations
Experimental results consistently confirm that better-calibrated propensity scores lead to superior performance. As illustrated in Tables II–V, methods utilizing calibrated propensity scores consistently outperform the raw baseline. This highlights a strong correlation between calibration quality and recommendation performance. Notably, Platt scaling achieves the highest calibration accuracy, which directly results in the best recommendation outcomes. In contrast, the raw baseline suffers from biased propensity scores that fail to accurately represent true interaction probabilities, leading to suboptimal recommendation results. Our experiments clearly show that this limitation can be effectively mitigated through proper calibration, underscoring the critical role of accurate propensity score estimation in achieving superior recommendation quality.
@@ IV-E Efficiency experiment
AUC DCG@K Recall@K
Baseline Dataset Methods K=2 K=4 K=6 K=2 K=4 K=6
Raw 0.7644$\pm$0.0032 0.7454$\pm$0.0158 1.0185$\pm$0.0196 1.2014$\pm$0.0126 0.8717$\pm$0.0214 1.4570$\pm$0.0243 1.9489$\pm$0.0209
Dropout 0.7673$\pm$0.0024 0.7496$\pm$0.0205 1.0271$\pm$0.0165 1.2073$\pm$0.0157 0.8835$\pm$0.0249 1.4764$\pm$0.0216 1.9603$\pm$0.0286
Ensembles 0.7805$\pm$0.0028 0.7546$\pm$0.0160 1.0254$\pm$0.0110 1.2132$\pm$0.0119 0.8835$\pm$0.0228 1.4637$\pm$0.0126 1.9684$\pm$0.0232
Dual FocalLoss 0.7699$\pm$0.0030 0.7542$\pm$0.0163 1.0106$\pm$0.0216 1.1939$\pm$0.0179 0.8834$\pm$0.0154 1.4651$\pm$0.0257 1.9514$\pm$0.0210
Coat Shopping Platt 0.7730$\pm$0.0018 0.7539$\pm$0.0201 1.0389$\pm$0.0208 1.2241$\pm$0.0170 0.8852$\pm$0.0235 1.4940$\pm$0.0290 1.9928$\pm$0.0208
Raw 0.7152$\pm$0.0053 0.5450$\pm$0.0093 0.7405$\pm$0.0097 0.8779$\pm$0.0091 0.6402$\pm$0.0113 1.0587$\pm$0.0143 1.4112$\pm$0.0141
Dropout 0.7130$\pm$0.0049 0.5501$\pm$0.0152 0.7597$\pm$0.0126 0.8774$\pm$0.0103 0.6539$\pm$0.0161 1.0822$\pm$0.0124 1.4183$\pm$0.0065
Ensembles 0.7258$\pm$0.0011 0.5431$\pm$0.0053 0.7491$\pm$0.0043 0.8747$\pm$0.0050 0.6510$\pm$0.0044 1.0923$\pm$0.0054 1.4293$\pm$0.0068
Dual FocalLoss 0.7121$\pm$0.0141 0.5358$\pm$0.0118 0.7447$\pm$0.0114 0.8788$\pm$0.0295 0.6429$\pm$0.0278 1.0584$\pm$0.0287 1.4186$\pm$0.0276
DR-JL Yahoo! R3 Platt 0.7248$\pm$0.0016 0.5532$\pm$0.0058 0.7555$\pm$0.0042 0.8816$\pm$0.0034 0.6602$\pm$0.0059 1.0926$\pm$0.0073 1.4314$\pm$0.0062
Raw 0.7691$\pm$0.0035 0.6830$\pm$0.0151 0.9661$\pm$0.0131 1.1578$\pm$0.0101 0.8185$\pm$0.0163 1.4261$\pm$0.0192 1.9409$\pm$0.0205
Dropout 0.7661$\pm$0.0011 0.7255$\pm$0.0200 0.9946$\pm$0.0138 1.1847$\pm$0.0127 0.8438$\pm$0.0257 1.4177$\pm$0.0247 1.9282$\pm$0.0312
Ensembles 0.7647$\pm$0.0038 0.7292$\pm$0.0273 1.0001$\pm$0.0192 1.1846$\pm$0.0182 0.8523$\pm$0.0352 1.4303$\pm$0.0190 1.9282$\pm$0.0299
Dual FocalLoss 0.7606$\pm$0.0052 0.7260$\pm$0.0205 1.0041$\pm$0.0162 1.1885$\pm$0.0203 0.8439$\pm$0.0222 1.4395$\pm$0.0171 1.9353$\pm$0.0225
Coat Shopping Platt 0.7728$\pm$0.0025 0.7648$\pm$0.0190 1.0155$\pm$0.0170 1.2223$\pm$0.0165 0.8987$\pm$0.0230 1.4688$\pm$0.0258 1.9957$\pm$0.0395
Raw 0.6678$\pm$0.0162 0.5371$\pm$0.0447 0.7441$\pm$0.0502 0.8636$\pm$0.0455 0.6383$\pm$0.0508 1.0808$\pm$0.0638 1.4006$\pm$0.0524
Dropout 0.6671$\pm$0.0135 0.5458$\pm$0.0194 0.7461$\pm$0.0240 0.8718$\pm$0.0225 0.6547$\pm$0.0209 1.0821$\pm$0.0363 1.4199$\pm$0.0382
Ensembles 0.6907$\pm$0.0026 0.5410$\pm$0.0093 0.7443$\pm$0.0098 0.8731$\pm$0.0090 0.6444$\pm$0.0120 1.0809$\pm$0.0133 1.4256$\pm$0.0127
Dual FocalLoss 0.6679$\pm$0.0461 0.5252$\pm$0.0062 0.7460$\pm$0.0194 0.8640$\pm$0.0203 0.6445$\pm$0.0191 1.0829$\pm$0.0166 1.4144$\pm$0.0190
MRDR Yahoo! R3 Platt 0.6988$\pm$0.0018 0.5623$\pm$0.0092 0.7571$\pm$0.0066 0.8858$\pm$0.0072 0.6687$\pm$0.0103 1.0862$\pm$0.0080 1.4326$\pm$0.009
TABLE IV: DRJL and MRDR CVR prediction performance on Coat Shopping and Yahoo! R3. The best results are shown in boldface and the second best results are marked using underline. ‡ indicates statistically significant improvements over the Raw method at p $<$ 0.05 level.
AUC DCG@K Recall@K
@@
Raw 0.6493$\pm$0.0034 0.4426$\pm$ 0.0076 0.6728$\pm$ 0.0110 0.8471$\pm$ 0.0108 0.5404$\pm$ 0.0096 1.0351$\pm$ 0.0166 1.5028$\pm$ 0.0163 0.8401
MC Dropout 0.6366$\pm$0.0058 0.4515$\pm$ 0.0063 0.6855$\pm$ 0.0064 0.8593$\pm$ 0.0069 0.5520$\pm$ 0.0073 1.0545$\pm$ 0.0080 1.5208$\pm$ 0.0101 0.8539
Deep Ensembles 0.6550$\pm$0.0019 0.4562$\pm$ 0.0045 0.6893$\pm$ 0.0048 0.8627$\pm$ 0.0042 0.5579$\pm$ 0.0052 1.0575$\pm$ 0.0063 1.5228$\pm$ 0.0058 0.8577
Dual FocalLoss 0.6373$\pm$0.0076 0.4634$\pm$0.0089 0.7068$\pm$0.0116 0.8705$\pm$0.0130 0.5648$\pm$0.0106 1.0761$\pm$0.0161 1.5420$\pm$0.0200 0.8706
IPS Platt Scaling 0.6682$\pm$0.0029 0.4657$\pm$ 0.0030 0.7021$\pm$ 0.0025 0.8774$\pm$ 0.0026 0.5677$\pm$ 0.0029 1.0754$\pm$ 0.0030 1.5455$\pm$ 0.0036 0.8723
Raw 0.6478$\pm$0.0022 0.4442$\pm$ 0.0083 0.6742$\pm$ 0.0111 0.8481$\pm$ 0.0115 0.5420$\pm$ 0.0096 1.0362$\pm$ 0.0160 1.5026$\pm$ 0.0170 0.8412
MC Dropout 0.6343$\pm$0.0074 0.4504$\pm$ 0.0042 0.6839$\pm$ 0.0044 0.8580$\pm$ 0.0052 0.5504$\pm$ 0.0041 1.0520$\pm$ 0.0047 1.5189$\pm$ 0.0071 0.8523
Deep Ensembles 0.6547$\pm$0.0030 0.4524$\pm$ 0.0064 0.6854$\pm$ 0.0039 0.8606$\pm$ 0.0043 0.5528$\pm$ 0.0051 1.0530$\pm$ 0.0048 1.5231$\pm$ 0.0067 0.8546
Dual FocalLoss 0.6436$\pm$0.0185 0.4673$\pm$0.0410 0.7065$\pm$0.0460 0.8815$\pm$0.0459 0.5698$\pm$0.0458 1.0834$\pm$0.0565 1.5530$\pm$0.0564 0.8769
DR-JL Platt Scaling 0.6679$\pm$0.0017 0.4701$\pm$ 0.0073 0.7070$\pm$ 0.0079 0.8834$\pm$ 0.0080 0.5733$\pm$ 0.0081 1.0823$\pm$ 0.0095 1.5553$\pm$ 0.0096 0.8786
Raw 0.5465$\pm$0.0064 0.4369$\pm$ 0.0246 0.6531$\pm$ 0.0287 0.8144$\pm$ 0.0295 0.5305$\pm$ 0.0292 0.9957$\pm$ 0.0376 1.4290$\pm$ 0.0401 0.8099
MC Dropout 0.5491$\pm$0.0044 0.4412$\pm$ 0.0125 0.6659$\pm$ 0.0183 0.8305$\pm$ 0.0200 0.5376$\pm$ 0.0152 1.0205$\pm$ 0.0275 1.4630$\pm$ 0.0302 0.8265
Deep Ensembles 0.5855$\pm$0.0050 0.4406$\pm$ 0.0101 0.6681$\pm$ 0.0127 0.8432$\pm$ 0.0135 0.5390$\pm$ 0.0131 1.0275$\pm$ 0.0171 1.4972$\pm$ 0.0215 0.8359
Dual FocalLoss 0.5813$\pm$0.0028 0.4777$\pm$0.0151 0.7165$\pm$0.0169 0.8920$\pm$0.0183 0.5828$\pm$0.0173 1.0955$\pm$0.0217 1.5661$\pm$0.0273 0.8884
MRDR Platt Scaling 0.6247$\pm$0.0035 0.4928$\pm$ 0.0057 0.7259$\pm$ 0.0076 0.8963$\pm$ 0.0079 0.5971$\pm$ 0.0070 1.0972$\pm$ 0.0109 1.5549$\pm$ 0.0119 0.8940
TABLE V: Overall performance on KuaiRand. ‡ indicates statistically significant improvements over the Raw method at p $<$0.05 level.
AUC DCG@K Recall@K
Datasets Methods K=2 K=4 K=6 K=2 K=4 K=6
MRDR-GPL 0.7521$\pm$0.0035 0.7488$\pm$ 0.0201 1.0061$\pm$ 0.0222 1.1949$\pm$ 0.0243 0.8734$\pm$ 0.0287 1.4219$\pm$ 0.0301 1.9283$\pm$ 0.0405
MRDR-CDR 0.7622$\pm$0.0031 0.7579$\pm$ 0.0201 1.0192$\pm$ 0.0192 1.1991$\pm$ 0.0198 0.8903$\pm$ 0.0204 1.4515$\pm$ 0.0277 1.9325$\pm$ 0.0402
DR-V2 0.7637$\pm$0.0039 0.7746$\pm$ 0.0188 1.0116$\pm$ 0.0164 1.2076$\pm$ 0.0176 0.8841$\pm$ 0.0184 1.4568$\pm$ 0.0152 1.9494$\pm$ 0.0324
ESCM2-DR 0.7681$\pm$0.0041 0.7528$\pm$ 0.0177 1.0273$\pm$ 0.0189 1.2081$\pm$ 0.0229 0.8945$\pm$ 0.0186 1.4810$\pm$ 0.0162 1.9662$\pm$ 0.0299
Coat Shopping MRDR-CAL(Ours) 0.7728$\pm$0.0025 0.7648$\pm$ 0.0190 1.0155$\pm$ 0.0170 1.2223$\pm$ 0.0165 0.8987$\pm$ 0.0230 1.4688$\pm$ 0.0258 1.9957$\pm$ 0.0395
MRDR-GPL 0.6617$\pm$0.0064 0.5384$\pm$ 0.0194 0.7369$\pm$ 0.0252 0.8605$\pm$ 0.0211 0.6408$\pm$ 0.0197 1.0657$\pm$ 0.0222 1.3982$\pm$ 0.0241
MRDR-CDR 0.6673$\pm$0.0035 0.5417$\pm$ 0.0162 0.7456$\pm$ 0.0123 0.8698$\pm$ 0.0125 0.6490$\pm$ 0.0202 1.0842$\pm$ 0.0182 1.4175$\pm$ 0.0147
DR-V2 0.6807$\pm$0.0026 0.5518$\pm$ 0.0125 0.7479$\pm$ 0.0143 0.8732$\pm$ 0.0156 0.658$\pm$ 0.0111 1.0784$\pm$ 0.0181 1.4154$\pm$ 0.0158
ESCM2-DR 0.6828$\pm$0.0161 0.5541$\pm$ 0.0144 0.7502$\pm$ 0.0126 0.8771$\pm$ 0.0171 0.6564$\pm$ 0.0156 1.0772$\pm$ 0.0175 1.4183$\pm$ 0.0154
Yahoo! R3 MRDR-CAL(Ours) 0.6988$\pm$0.0018 0.5623$\pm$ 0.0092 0.7571$\pm$ 0.0066 0.8858$\pm$ 0.0072 0.6687$\pm$ 0.0103 1.0862$\pm$ 0.0080 1.4326$\pm$ 0.0090
MRDR-GPL 0.5401$\pm$0.0084 0.4335$\pm$ 0.0123 0.6446$\pm$ 0.0178 0.8049$\pm$ 0.0143 0.5261$\pm$ 0.0171 0.9792$\pm$ 0.0178 1.4101$\pm$ 0.0146
MRDR-CDR 0.5498$\pm$0.0054 0.4326$\pm$ 0.0098 0.6573$\pm$ 0.0145 0.8297$\pm$ 0.0132 0.5289$\pm$ 0.0126 1.0111$\pm$ 0.0146 1.4739$\pm$ 0.0165
DR-V2 0.5765$\pm$0.0047 0.4465$\pm$ 0.0078 0.6795$\pm$ 0.0121 0.8533$\pm$ 0.0098 0.5452$\pm$ 0.0100 1.0459$\pm$ 0.0122 1.5117$\pm$ 0.0123
ESCM2-DR 0.5836$\pm$0.0068 0.4773$\pm$ 0.0101 0.7059$\pm$ 0.0155 0.8760$\pm$ 0.0121 0.5781$\pm$ 0.0143 1.0694$\pm$ 0.0143 1.5258$\pm$ 0.0152
KuaiRand MRDR-CAL(Ours) 0.6247$\pm$0.0035 0.4928$\pm$ 0.0057 0.7259$\pm$ 0.0076 0.8963$\pm$ 0.0079 0.5971$\pm$ 0.0070 1.0972$\pm$ 0.0109 1.5549$\pm$ 0.0119
TABLE VI: The comparison with the SOTA methods, ‡ indicates statistically significant improvements over the ESCM2-DR method at p $<$ 0.05 level.